Express the wavelength of 500 nm (green light) in units of wavenumbers (c work nt: The wavenumber (cm1) is a common unit used in infrared spectroscopy number of wavelengths that would fit in 1 cm. 2 x 104 cm1 2 x 106 cm 2 x 102 cm1 -1 2 cm
Answer #1

given,
wavelength = 500 nm
= (500*10^-7) cm
= (5*10^-5) cm

use,
wavenumber = 1/(wavelength)
= 1/(5*10^-5)
= (0.2*10^5) cm^-1
= (2*10^4) cm^-1

Answer: (2*10^4) cm^-1